OpenStudy (anonymous):

solve each system by elimination. 3x+y=-26 2x-y=-19 what would be X and Y

OpenStudy (anonymous):

first, change the value of y 2x-y=-19 y=2x+19

OpenStudy (anonymous):

second, replace y in 3x+y=-26 3x+(2x+19)=-26

OpenStudy (jdoe0001):

\( \begin{array}{llll} 3x+y=-26\\ 2x-y=-19\\ \hline\\ 5x+0y=-45 \end{array}\bf\implies 5x=-45\) solve for "x", what do you get?

OpenStudy (anonymous):

for x you would get -9 but i can't get something for y that works for both equations

OpenStudy (anonymous):

it will find x, then replace the value of x you have got to 2x-y=-19 you will have y happy to help...

OpenStudy (jdoe0001):

well, now that we know that x = -9, plug that in either equation, say the 1st one, so \(\bf 3x+y=-26\implies 3(-9)+y=-26\implies y=?\)

OpenStudy (anonymous):

Don't you -27 to -26 an then that would be y

OpenStudy (anonymous):

i mean add -27 to -26 and then that would be y

OpenStudy (jdoe0001):

\(\bf 3x+y=-26\implies 3(-9)+y=-26\implies -27+y=-26\\ \quad \\ \cancel{+27-27}+y=-26+27\)

OpenStudy (anonymous):

so y would be 1 right

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

\(\begin{array}{llll} 3x+y=-26\\ 2x-y=-19\\ \hline\\ 5x+0y=-45 \end{array}\bf\implies 5x=-45\) notice, we already has a +y and -y right on top on one another, so we just "elimiinate" it, thus

OpenStudy (anonymous):

okay i see that know thank you very much that helped a lot

OpenStudy (jdoe0001):

yw